i am having a difficult time trying to solve:
6x - 9y= 25.5
7y - 3x= -14
i cannot figure out what i need to do to get the x's and y's to line up in order to solve the equation...

Answer by InfantileBear(2)

You start with the two original equations:
6x-9y=25.5
7y-3x=-14
Take the second equation and multiply it by 2 to get:
14y-6x=-28
Now you can add the first equation to this equation to cancel out the x's:
(14y-6x=-28)
+(-9y+6x=25.5)
This will give you 5y=-2.5
If you solve for y, you get y=-.5 and if you plug this number back into one of the original equations,:
6x-9(-.5)=25.5
you should have x=3.5
And to check your answer, plug the x and y that you got into one of the original equations and see if it works out.